Velocity of a particle is given by v= 3-t m/s. The distance travelled by particle in 1st four second is

Grade:

2 Answers

sayantani bardhan
6 years ago
at t=0 , v = 3-0= 3 m/s v=u= 3 m/sat t=4 , v= 3 - 4 = -1m/snow accleration(a) = (v-u)/tso a = (-1-3)/4 = -1m/s^2now putting the values in the equation v^2 = u^2 + 2 we gets = 4 m
Ashish kumar behra
5 years ago
Integrate the equation ,v=3-t After intigrating We get s=3t-(t^2/2)Now puting 4 in place of t we get thenWe get the distance equal to 4meterS=4m